OpenStudy (fanduekisses):

The half-life of a certain radioactive substance is 80 days. There are 4 grams initially. When will there be 1 gram remaining?

OpenStudy (fanduekisses):

Use the standard formula \[A(t)= A_{0}e^{rt}\]

OpenStudy (fanduekisses):

Is this the right set up? \[1= 4e^{80t}\]

OpenStudy (fanduekisses):

the rate of decay is 80?

OpenStudy (freckles):

how did you get that rate ?

OpenStudy (freckles):

We are given by first sentence the following: \[\frac{1}{2} A_0=A_0e^{r \cdot 80} \\ \text{ solve for } r \\ \text{ where } r \text{ is the rate } \\ t=80 \text{ is the amount of days \it takes for some amount to get half }\]

OpenStudy (freckles):

once you find r you can back to that 1=4*e^(rt) where you know r now from the previous equation I just asked you to solve

OpenStudy (fanduekisses):

is r= ln(2/4) / 80 ?

OpenStudy (freckles):

2/4 is the same as 1/2

OpenStudy (fanduekisses):

oh ok just reduced

OpenStudy (freckles):

so you can write as that or this last line I have here \[r=\frac{1}{80} \ln(\frac{1}{2}) \\ r=\frac{1}{80}(\ln(1)-\ln(2)) \\ r=\frac{1}{80}(0-\ln(2)) \\ r=\frac{-\ln(2)}{80}\]

OpenStudy (freckles):

\[1=4 e^{ rt } \\ \text{ where we found } r=\frac{-\ln(2)}{80} \\ 1=4 e^{\frac{-\ln(2)}{80} t} \\ \text{ solve for } t \]

OpenStudy (freckles):

mention me if you have trouble my notifications aren't workiong

OpenStudy (fanduekisses):

@freckles hey I get t= 160 :}

OpenStudy (freckles):

one sec checking...

OpenStudy (freckles):

sounds great